OpenStudy (anonymous):

Find the derivative of the following function:

OpenStudy (anonymous):

Oh good heavens just look at the time!

OpenStudy (anonymous):

\[\frac{ 2(\sqrt[5]{x})-\sqrt[7]{x^4} }{ x^2 +3x + 1 }\]

OpenStudy (idku):

you can use do the quotient rule, re-writing the rules as powers. That can be a shi(t)load of work though...

OpenStudy (idku):

I mean re-writing the roots....

OpenStudy (anonymous):

Yeah I meant to add how to do the quotient rule. I never really understood.

OpenStudy (anonymous):

I could just give an easier question if you like

OpenStudy (idku):

can you re-write the functions, so that numerator will have powers of x, instead of roots of x?

OpenStudy (anonymous):

Sure

OpenStudy (idku):

go ahead:)

OpenStudy (anonymous):

\[\frac{ 2(x^\frac{ 1 }{ 5 }) -x^\frac{ 4 }{ 7 } }{ x^2 +3x + 1 }\]

OpenStudy (idku):

don't worry about putting fractions like that, for future reference use ^{1/5} for example... yes that is right.

OpenStudy (idku):

what is the derivative of the denominator? what is the derivative of the numerator?

OpenStudy (anonymous):

denominator is (x^2 +3x +1)^2 What would be the numerator though. This is where I struggle the most

OpenStudy (idku):

\[quotient:~~~\frac{d}{dx}~\frac{f(x)}{g(x)}=~\frac{g(x)f'(x)~-~f(x)g'(x)}{[~g(x)~]^2}\]

OpenStudy (idku):

so you would need the derivative of numerator, and of denominator to find out what the top is going to be.

OpenStudy (idku):

f'(x) & g'(x) is a derivative of a function notation, you know that, right?

OpenStudy (anonymous):

sorry let me write this down, you've already written something I might have missed in class.

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Okay, ill try and finish the question with the formula.

OpenStudy (idku):

sure

OpenStudy (anonymous):

\[\frac{ (\frac{ 2 }{ 5 }x^{-4/5} - \frac{ 4 }{ 7 }x^{-3/7})(x^2 +3x +1) - 2x+3(2x^{1/5} - x^{4/7}) }{ (x^2+3x+1)^2 }\]

OpenStudy (anonymous):

that took a while haha

OpenStudy (idku):

a little mistake,.

OpenStudy (idku):

-(2x+3)(2x....)

OpenStudy (anonymous):

oh yeah I missed that, its done the right way on my paper

OpenStudy (idku):

very good!

OpenStudy (anonymous):

haha thanks! I really appreciate your help! :)

OpenStudy (idku):

yw